# From Samples to Scenarios: A New Paradigm for Probabilistic Forecasting

### v1

@@ A.2.3 Optimal Probabilities via Probability Loss
Proposition 2. The optimal probability vector $\mathbf{p}$ produced by TimePrism converges to the true conditional probability mass function over the Voronoi regions. That is, $p_{n}\to P(\mathbf{y}_{\text{gt}}\in R_{n}\mid\mathbf{x})$, where $R_{n}$ is the Voronoi region of the $n$-th scenario.
Proof. The optimization objective for the probability loss is to minimize the expected Cross-Entropy loss:
$\mathbb{E}[\mathcal{L}_{\text{prob}}]=\mathbb{E}_{\mathbf{x},\mathbf{y}_{\text{gt}}}[\text{CrossEntropy}(\bm{\pi}(\mathbf{x}),n^{*}(\mathbf{x},\mathbf{y}_{\text{gt}}))]$ (19)
Let $q(n\mid\mathbf{x})=P(\mathbf{y}_{\text{gt}}\in R_{n}\mid\mathbf{x})$ be the true, unknown probability that the $n$-th scenario is the winner for a given history $\mathbf{x}$. The expected loss is then the cross-entropy between the true distribution $\mathbf{q}(\mathbf{x})$ and the model’s predicted distribution $\mathbf{p}(\mathbf{x})=\text{Softmax}(\bm{\pi}(\mathbf{x}))$, averaged over all histories $\mathbf{x}$. By the properties of cross-entropy, this is equivalent to minimizing the Kullback-Leibler (KL) divergence plus the entropy of the true distribution:
@@ A.2.4 Probability Matching and CRPS
Our paradigm’s ability to achieve strong performance on the Weighted CRPS metric is rooted in its direct optimization of a true probability distribution. As established in Proposition 2, the Cross-Entropy loss drives the model’s output probability vector, $\mathbf{p}=\text{Softmax}(\bm{\pi})$, to match the true conditional probability mass function over the set of optimal scenarios. The objective is to minimize the Kullback-Leibler (KL) divergence between the predicted and true discrete distributions, $D_{KL}(\mathbf{q}(\mathbf{x})\|\mathbf{p}(\mathbf{x}))$, where $\mathbf{q}(\mathbf{x})$ is the true distribution of winner outcomes. Since the Weighted CRPS directly incorporates the probability vector $\mathbf{p}$ (Eq. equation 4), a model that learns a more accurate probability distribution is expected to achieve a lower (better) score.
@@ D.1 Results of Complementary Metrics
D.2 Experiments on TimePrism Configuration
D.2.1 History Length
@@
For a more direct comparison, we present an additional experiment in Table 7 where TimePrism uses the full available history, a variant we term "Full History" (Full His.). The length of this history is set to be comparable to the total context available to the baselines’ data processing modules as in (Cortés et al., 2025). The results show that using a longer history does not consistently improve TimePrism’s performance; in some cases, the scores are similar or slightly worse, though still highly competitive. This is not a perfectly fair comparison, as other models are designed with feature engineering capabilities to extract value from long lagged inputs, while our simple linear model uses the full history directly. For such a simple architecture, a much longer input sequence can introduce noise without a sophisticated mechanism to filter it, which explains why more data does not necessarily lead to better performance.
@@
D.2.2 Number of Scenarios
The number of scenarios, $N$, is the primary hyperparameter of TimePrism. We conduct an ablation study on the Solar dataset to analyze its impact on performance, with results shown in Table 8. As $N$ increases, the performance on our primary metrics, CRPS and Distortion, generally improves. This demonstrates that a larger set of scenarios allows the model to better represent the underlying probability space. Notably, even in its deterministic form ($N=1$), which is equivalent to a DLinear-like architecture, TimePrism provides a highly competitive baseline, outperforming several complex probabilistic models on this dataset. This highlights the strength of the underlying decomposition-based structure. However, the more significant performance gains are achieved when the model is allowed to produce multiple scenarios, confirming the effectiveness of the Probabilistic Scenarios paradigm.
While the primary metrics continue to improve up to $N=1024$, the performance gain from 625 to 1024 is marginal, and the MSE score even slightly increases. As our primary goal is to validate the paradigm with a simple and efficient configuration rather than to exhaustively optimize the model, we selected $N=625$ as a balanced and representative setting for our main experiments.
Table 8: Effect of the Number of Scenarios (NN) on Performance. All metrics are evaluated on the Solar dataset. The best result in each column is in bold, and the second best is underlined.
Scenarios CRPS Distortion MAE MSE
1 0.202 0.270 0.254 0.077
16 0.135 0.260 0.216 0.060
256 0.094 0.149 0.105 0.068
625 0.084 0.140 0.098 0.062
1024 0.083 0.084 0.078 0.082
D.3 Additional Visualization and Qualitative Analysis
@@
D.3.1 Visualizations on Other Datasets
@@
D.3.2 Full Comparison on Solar

### v2

@@ A.2.3 Optimal Probabilities via Probability Loss
Proposition 2. At the global minimum of the expected probability loss, the predicted probability vector $\mathbf{p}$ matches the true conditional probability mass function over the Voronoi regions. That is, $p_{n}=P(\mathbf{y}_{\text{gt}}\in R_{n}\mid\mathbf{x})$, where $R_{n}$ is the Voronoi region of the $n$-th scenario.
Proof. The optimization objective for the probability loss is to minimize the expected Cross-Entropy loss. We denote the cross-entropy between two discrete distributions $\mathbf{q}$ and $\mathbf{p}$ as $H(\mathbf{q},\mathbf{p})=-\sum_{n}q_{n}\log p_{n}$.
$\mathbb{E}[\mathcal{L}_{\text{prob}}]=\mathbb{E}_{\mathbf{x}}\left[\mathbb{E}_{\mathbf{y}_{\text{gt}}|\mathbf{x}}[-\log p_{n^{*}(\mathbf{x},\mathbf{y}_{\text{gt}})}(\mathbf{x})]\right]=\mathbb{E}_{\mathbf{x}}[H(\mathbf{q}(\mathbf{x}),\mathbf{p}(\mathbf{x}))]$ (19)
Let $q(n\mid\mathbf{x})=P(\mathbf{y}_{\text{gt}}\in R_{n}\mid\mathbf{x})$ be the true, unknown probability that the $n$-th scenario is the winner for a given history $\mathbf{x}$. The inner expectation corresponds to the cross-entropy between this true distribution $\mathbf{q}(\mathbf{x})$ and the model’s predicted distribution $\mathbf{p}(\mathbf{x})=\text{Softmax}(\bm{\pi}(\mathbf{x}))$. By the properties of cross-entropy:
@@ A.2.4 Probability Matching and CRPS
Our paradigm’s ability to achieve strong performance on the Weighted CRPS metric is rooted in its direct optimization of a true probability distribution. As established in Proposition 2, the Cross-Entropy loss drives the model’s output probability vector, $\mathbf{p}=\text{Softmax}(\bm{\pi})$, to match the true conditional probability mass function over the set of optimal scenarios. The objective is to minimize the Kullback-Leibler (KL) divergence between the predicted and true discrete distributions, $D_{KL}(\mathbf{q}(\mathbf{x})\|\mathbf{p}(\mathbf{x}))$, where $\mathbf{q}(\mathbf{x})$ is the true distribution of winner outcomes. Since the Weighted CRPS directly incorporates the probability vector $\mathbf{p}$ (equation 4), a model that learns a more accurate probability distribution is expected to achieve a lower (better) score.
@@
A.3 Connection to Discrete Representation Learning
As noted in our discussion on the model’s effectiveness, the Probabilistic Scenarios paradigm shares conceptual roots with discrete representation learning techniques, most notably Vector Quantized Variational AutoEncoders (VQ-VAE) (van den Oord et al., 2017). Both approaches posit that continuous spaces can be effectively approximated by a finite set of discrete vectors. However, TimePrism distinguishes itself from VQ-VAE in three fundamental aspects, tailored for the forecasting task:
  • Discretization Target: VQ-VAE discretizes latent features, which serve as intermediate representations. In contrast, TimePrism directly discretizes future trajectories, operating within the final output space.
  • Nature of Codebook/Scenarios: VQ-VAE utilizes a static, global codebook shared across all inputs, where codes are fixed parameters learned from the entire dataset. Conversely, TimePrism generates a dynamic set of scenarios in real-time based on the input. These scenarios function as a conditional codebook that adapts to the specific history of each time series.
  • Probability Modeling: VQ-VAE typically employs an implicit, two-stage approach that requires training a separate prior model over discrete codes to perform sampling and probability estimation. TimePrism, however, uses an explicit, end-to-end approach featuring a built-in probability head that directly outputs the probability distribution $p$ over the generated scenarios in a single forward pass.
@@ D.1 Results of Complementary Metrics
D.2 Experiments on History Length Configuration
@@
For a more direct comparison, we present an additional experiment in Table 8 where TimePrism uses the full available history, a variant we term "Full History" (Full His.). The length of this history is set to be comparable to the total context available to the baselines’ data processing modules as in Cortés et al. (2025). The results show that using a longer history does not consistently improve TimePrism’s performance; in some cases, the scores are similar or slightly worse, though still highly competitive. This is not a perfectly fair comparison, as other models are designed with feature engineering capabilities to extract value from long lagged inputs, while our simple linear model uses the full history directly. For such a simple architecture, a much longer input sequence can introduce noise without a sophisticated mechanism to filter it, which explains why more data does not necessarily lead to better performance.
@@
D.3 Paradigm Generalizability: Adapting to Transformer Architectures
Table 9: Generalizability Analysis with Transformer Backbone. Comparison of Primary Metrics (CRPS and Distortion) across five datasets. TimePrism-iT represents the iTransformer (Liu et al., 2023) architecture adapted to our Probabilistic Scenarios paradigm. All experiments use Seed 3141.
Model Elec. Exch. Sol. Traf. Wiki.
CRPS Dis. CRPS Dis. CRPS Dis. CRPS Dis. CRPS Dis.
DeepAR 0.993 2.79 0.698 1.89 0.607 1.11 0.829 1.82 1.41 4.88
TimeGrad 0.230 0.720 0.739 1.33 0.237 0.587 0.163 0.540 0.516 1.62
TempFlow 0.449 1.73 0.988 1.55 0.278 0.555 0.613 1.01 1.81 71.6
Trf.Flow 0.449 1.73 0.988 1.55 0.278 0.555 0.613 1.01 1.81 71.6
Tactis2 0.285 0.637 0.641 0.919 0.222 0.567 0.243 0.55 0.481 1.37
TimeMCL 0.375 0.603 1.30 1.10 0.301 0.485 0.251 0.455 0.624 1.32
TimePrism 0.148 0.237 0.456 0.588 0.0835 0.140 0.109 0.140 0.508 1.01
TimePrism-iT 0.330 0.600 0.454 0.681 0.164 0.245 0.201 0.371 0.756 1.425
To more rigorously validate that the superior performance of our method stems from the proposed Probabilistic Scenarios paradigm rather than solely the specific linear architecture of TimePrism, we conducted a controlled study adapting a distinct, complex architecture to our framework. We selected one of the state-of-the-art Transformer-based time series models, iTransformer (Liu et al., 2023), as the backbone.
Experimental Setup. We developed a variant named TimePrism-iT, where the linear encoder of TimePrism is replaced by the inverted Transformer structure from Liu et al. (2023). Crucially, to demonstrate the "out-of-the-box" applicability and robustness of our paradigm, we did not perform extensive hyperparameter tuning for TimePrism-iT. Instead, we applied a generally consistent configuration across all datasets. This setup serves as a rigorous stress test to verify if the paradigm can yield performance gains without relying on architecture-specific optimization.
Results Analysis. The comparative results are presented in Table 9. Despite being an unoptimized implementation, TimePrism-iT demonstrates remarkable performance. It outperforms standard baselines on 6 out of 10 metrics across the five datasets. Notably, on the Exchange dataset, TimePrism-iT achieves a CRPS of 0.454, slightly surpassing even the original linear TimePrism (0.456).
D.4 Extended Evaluation on Additional Benchmarks
Table 10: The results of models in datasets from GIFT-Eval (Aksu et al., 2024) and fev-bench Shchur et al. (2025). The best result in each column is in bold.
Dataset UCI Hosp. Hier. M-Den.
Metrics CRPS Distortion CRPS Distortion CRPS Distortion CRPS Distortion
ETS 0.450 0.811 0.585 1.30 0.990 4.03 0.782 2.23
Tactis2 0.605 0.787 0.583 1.20 0.623 1.58 0.614 1.13
TimeMCL 0.359 0.449 0.722 1.13 1.08 1.33 0.771 0.597
TimePrism 0.261 0.394 0.565 1.06 0.602 1.03 0.907 1.11
To provide a more comprehensive evaluation of our proposed paradigm, we extended our experiments to include four datasets selected from two latest benchmarks: Gift-Eval (Aksu et al., 2024) and fev-bench (Shchur et al., 2025). These datasets were chosen to cover diverse domains: Hierarchical Sales (Retail, abbr. Hier.), M-DENSE (Mobility, abbr. M-Den.), Hospital Admissions (Healthcare, abbr. Hosp.), and UCI Air Quality (Nature, abbr. UCI).
Experimental Setup. For these experiments, we selected the numerical baseline ETS and the two most competitive neural models from Table 1 and Table 2, namely TimeMCL and Tactis2, for comparison. All models were evaluated using a random seed of 3141 to ensure reproducibility.
Results Analysis. As shown in the additional results in Table 10, TimePrism maintains its strong performance across these new domains. Regarding the M-DENSE dataset, we observed that TimePrism exhibits relatively higher distortion. We hypothesize that the nature of this dataset may be more suitable for RNN backbones, as both Tactis-2 and TimePrism perform suboptimally on this dataset, while TimeMCL remains competitive. This is not a limitation of our new paradigm, but rather a consequence of TimePrism’s simple structure. Nevertheless, achieving SOTA results in 15 out of 18 metrics across 9 datasets still demonstrates the effectiveness of the TimePrism model and highlights the potential of the new paradigm.
D.5 Probability Calibration Diagnostics
To rigorously assess the reliability of the probabilities assigned by TimePrism, we employ two standard diagnostic tools: the Reliability Diagram (Coverage vs. Nominal Confidence) and the Probability Integral Transform (PIT) Histogram.
Methodology. Since TimePrism outputs a tuple $(\mathcal{Y}_{\text{pred}},\mathbf{p})$ consisting of a finite set of scenarios $\mathcal{Y}_{\text{pred}}=\{\mathbf{y}_{n}\}_{n=1}^{N}$ and their associated probabilities $\mathbf{p}=(p_{1},\dots,p_{N})$, we compute these metrics as follows:
  • PIT Histogram: For a ground truth observation $\mathbf{y}_{\text{gt}}$, the PIT value is the cumulative probability of scenarios that are less than or equal to the observation: $\text{PIT}=\sum_{n=1}^{N}p_{n}\cdot\mathbb{I}(\mathbf{y}_{n}\leq\mathbf{y}_{\text{gt}})$. For a perfectly calibrated model, the distribution of PIT values over the test set should approach a Uniform distribution $U[0,1]$, resulting in a flat histogram.
  • Reliability Diagram: We calculate the empirical coverage for varying nominal confidence levels $\alpha\in[0,1]$. The prediction interval for a level $\alpha$ is constructed by aggregating the scenarios $\mathbf{y}_{n}$ with the highest probabilities until their cumulative sum reaches $\alpha$. If the model is well-calibrated, the curve should align with the diagonal $y=x$. Curves above the diagonal indicate under-confidence (conservative), while curves below indicate over-confidence.
(a) PIT (Exchange)
(b) Reliability (Exchange)
(c) PIT (Solar)
(d) Reliability (Solar)
Figure 5: Calibration Diagnostics. The diagnostics show different behaviors across datasets: (a)(b) Exchange dataset demonstrates near-perfect calibration; (c)(d) Solar dataset exhibits a slightly conservative profile to ensure robust tail coverage.
Analysis. We performed these diagnostics on two representative datasets: Exchange and Solar. The results are visualized in Figure 5.
  • Exchange Dataset (Fig. 5(a) & 5(b)): The diagnostics indicate near-perfect calibration. The PIT histogram is remarkably flat, and the Reliability Diagram closely follows the ideal diagonal line. This suggests that for stable financial data, TimePrism accurately estimates the true uncertainty distribution.
  • Solar Dataset (Fig. 5(c) & 5(d)): The diagnostics exhibit a slightly conservative profile. The PIT histogram shows a mild hump shape, and the Reliability curve lies slightly above the diagonal. This behavior is expected and often desirable for highly stochastic, multimodal data like Solar energy. It indicates that TimePrism tends to widen its predicted scenario distribution to safely encompass multimodality and potential outliers. This "conservative" strategy ensures robust coverage of low-probability, high-impact tail events without becoming over-confident, aligning with our design goal of prioritizing coverage adequacy.
D.6 Additional Visualization and Qualitative Analysis
@@
D.6.1 Visualizations on Other Datasets
@@
D.6.2 Full Comparison on Solar